In laying out an article of 32,000 words, an editor decides to use two sizes of type. Using the larger type, a printed page contains 900 words. Using the smaller type, a page contains 1,400 words. The article is allotted 30 full pages in a magazine. How many pages must be in the smaller type?
Answer by stanbon(75887)

In laying out an article of 32,000 words, an editor decides to use two sizes of type. Using the larger type, a printed page contains 900 words. Using the smaller type, a page contains 1,400 words. The article is allotted 30 full pages in a magazine. How many pages must be in the smaller type?
----
Pages Equation: n + f = 30 pages
----
Wods Equation:900n + 1400f = 32000
----------------------------------
Modify both equations:
9n + 14f = 320
9n + 9f = 9*30
----
Subtract and solve for "f":
5f = 50
f = 10 (number of 1400 word pages)
-----
Solve for "n":
n + f = 30
n + 10 = 30
n = 20 pages (number of 900 word pages)
============================================
Cheers,
Stan H.